OpenStudy (darkbluechocobo):

Find the x-intercepts of the following quadratic function. f(x) = -x^2 + 9

OpenStudy (darkbluechocobo):

i got (3,0) and (-3,0)

OpenStudy (anonymous):

3 and -3 is correct.

OpenStudy (cwrw238):

write it as 9 - x^2 this is a special type of qudratic called the difference of 2 squares take square root of both terms and insert + in one bracket and - in the other 9 - x^2 = (3 - x)(3 + x) = 0 so x intercepts are 3 , -3

OpenStudy (darkbluechocobo):

I like this one a lot better than the quadratic ones :p a lot simpler

ganeshie8 (ganeshie8):

me too :) Kindly show your work so that others know what you need help with exactly you may use wolfram for checking your answers http://www.wolframalpha.com/input/?i=x+intercepts+-x%5E2+%2B+9